OpenStudy (anonymous):

(2)((-3squarert40)/49)) how do you simplify???

OpenStudy (anonymous):

\[2(-3\sqrt{40}/49)\]

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

i would first multiply by 49 to get rid of the fraction..

OpenStudy (tkhunny):

Please don't multiply this expression by anything. That would be a different expression. Simplify \(\sqrt{40}\)

OpenStudy (anonymous):

why would he not multiply by 49 so that way its just sqrt40?

OpenStudy (anonymous):

@tkhunny why did you just leave

OpenStudy (tkhunny):

You would not multiply by 49 because that is an entirely different expression. Would you even think of simplifying 2/3 by multiplying by 3?

OpenStudy (anonymous):

:) thanks yall but what would i do next??\[(2)(-6\sqrt{10}/49)\]

OpenStudy (tkhunny):

You're about done. 2(-6) = -12

OpenStudy (anonymous):

:) thanks!!